Alden Company uses a three-variance analysis for factory overhead variances. Practical capacity is defined as 34 setups and 34,000 machine hours to manufacture 8,500 units for the year. Selected data for 2019 follow: Budgeted fixed factory overhead: Setup cost $ 85,000 Other 186,000 $ 271,000 Total factory overhead cost incurred $ 485,000 Variable factory overhead rate: Per setup $ 400 Per machine hour $ 5.00 Total standard machine hours allowed for the units manufactured 27,000 hours Machine hours actually worked 32,500 hours Actual total number of setups 30 Actual number of units produced during the year 6,750 Standard number of setups for units produced during the year 27 Required: 1. Compute: (a) the total overhead spending variance, (b) the overhead efficiency variance, and (c) the total overhead flexible-budget variance for 2019. Label each variance as favorable (F) or unfavorable (U). 2. Assume that the company includes all setup costs as variable factory overhead. The budgeted total fixed overhead, therefore, is $186,000, and the standard variable overhead rate per setup is $2,900. What is: (a) the total overhead spending variance, (b) the overhead efficiency variance, and (c) the total overhead flexible-budget variance for the year? Label each variance as favorable (F) or unfavorable (U). 3. Assume that the company uses only machine hours as the activity measure to apply both variable and fixed overhead, and that it includes all setup costs as variable factory overhead. What is (a) the Total Overhead Spending Variance, (b) the Overhead Efficiency Variance, and (c) total Overhead Flexible-Budget Variance for the year? Indicate whether each variance is favorable (F) or unfavorable (U).
Answer #1

Requirement #1 )

Standard Manufacturing Hours(MH) per unit=34000/8500=4hrs/ unit

Standard no. of units manufactured:

Standard MH allowed for the units manufactured/Standard MH per unit

=27,000/4

=6,750units

Budgeted no. of units per set up=Practical production/Practical setups

=8500/34

=250 units

Standard no of setups for actual production =Actual pdn/250

=6750/250

=27 setups

a) Spending variance =(Budgeted VOH+FOH)°° - Actual OH

=(((30*400)+(5*32500)+271000))-485000

=445500-485000

=39500 Unfavorable

°°For actual output

b) Efficiency variance =(Budgeted VOH+FOH for budgeted output) - (Budgeted VOH+FOH for actual output  

​​​​​​=((27*400)+(5*27000)+271000)-445500

=416800-445500

=28700 Unfavorable

c) ​​​​​​Total OH Flexible budget variance =Spending variance + Efficiency variance

=39500+28700

=68200 Unfavorable

Requirement #2)

a) Spending variance =((30*2900)+(5*32500)+186000)-485000

=435500-485000

=49500 Unfavorable

b) Efficiency variance =((27*2900)-(5*27000)+186000)-435500

=399300-435500

=36200 Unfavorable

c) Total OH flexible budget variance =49500+36200

=85700 Unfavorable

AS PER HOMEWORKLIB RULES I'VE ANSWERED THE FIRST 2 PARTS AND TOTAL 6 SUB-PARTS. PLEASE POST THE SAME QUESTION ONCE AGAIN ASKING TO SOLVE THE REMAINING PORTION...PLEASE DON'T GIVE ME A NEGATIVE RATING FOR THIS REASON.IN CASE OF ANY DOUBTS FEEL FREE TO COMMENT BELOW